Eveeeeee2007:

A pole that is 3.5 m tall casts a shadow that is 1.37 m long. At the same time, a nearby building casts a shadow that is 46.25m long. How tall is the building?

Vocaloid:

The height of an object is proportional to its shadow So (height of the pole) / (length of shadow of pole) = (height of building) / (shadow of building) we want to know the actual building's height, so we let x = height of building. plugging in x, along w/ the other numbers: 3.5 / 1.37 = x / 46.25 solve for x